# Imputation-free Learning of Tabular Data with Missing Values using Incremental Feature Partitions in Transformer

### v4

@@ Abstract
Tabular data sets with varying missing values are prepared for machine learning using an arbitrary imputation strategy. Synthetic values generated by imputation models often raise concerns about data quality and the reliability of data-driven outcomes. To address these concerns, this article proposes an imputation-free incremental attention learning (IFIAL) method for tabular data. A pair of attention masks is derived and retrofitted to a transformer to directly streamline tabular data without imputing or initializing missing values. The proposed method incrementally learns partitions of overlapping and fixed-size feature sets to enhance the efficiency and performance of the transformer. The average classification performance rank order across 17 diverse tabular data sets highlights the superiority of IFIAL over 11 state-of-the-art learning methods with or without missing value imputations. Further experiments substantiate the robustness of IFIAL against varying missing value types and rates compared to methods involving missing value imputation. Our analysis reveals that a feature partition size of half the original feature space is, both computationally and in terms of accuracy, the best choice for the proposed incremental learning. The proposed method is one of the first solutions to enable deep attention learning of tabular data without requiring missing-value imputation. The source code for this paper is available
@@ 1 Introduction
Missing values at varying rates and types are ubiquitous in health science involving electronic health records (EHR) data. For two reasons, missing values remain critical to ensuring unbiased data and reliable data-driven outcomes. First, the state-of-the-art machine learning (ML) methods that promise precision medicine are not designed for data with missing values. Second, missing values are handled arbitrarily by excluding samples and features or imputing with model-generated synthetic values to address the first reason. Sample exclusion biases medical data with sicker patients when their healthy counterparts are known to have more missing values [1]. Feature exclusion due to high missing rates can miss potentially important predictors. In contrast, many health scientists remain skeptical of using medical data imputed with model-generated synthetic values, especially using a hard-to-interpret ML model. Synthetic values generated by models can inadvertently alter feature importance, data statistics, and correlations between features, dampening the predictive ability of data [2]. Multiple imputations using chained equations (MICE) is arguably the most popular method in health science to impute missing values. MICE estimates missing values using a chain of linear regression models for individual features with two major limitations. First, missing values of one feature may be independent of other features, making regression models ineffective. Second, the relationship between features is rarely linear in practice. In other words, data statistics and missing value patterns should guide the modeling and selection of an imputation algorithm. Unfortunately, there is no best imputation method for all data sets or missing value types [3, 4, 5]. All concerns surrounding imputation methods underscore the need for alternatives to imputation in learning data with missing values. This paper proposes No Imputation Incremental Learning (NIIL) of tabular data with missing values. NIIL can directly classify tabular data with missing values, eliminating biases and concerns surrounding data exclusions and synthetic value imputation.
@@ 1.1 Related work
Feature incremental learning (FIL) is one of the cornerstones of deep learning (DL), which is not viable in traditional ML. Although highly successful with homogeneous image features [6], incremental learning is challenging in the heterogeneous feature space of tabular data [7]. Moreover, the superiority of traditional ML in tabular data over deep learning is well documented in the literature [8]. Some recent work has achieved FIL on tabular data [7, 9, 10, 11]. A common approach to FIL trains multiple models separately to accommodate new features and aggregate model predictions through a weighting mechanism [10, 7]. Alternatively, mamba [9] and the transformer-based method [11] use a single model to train on additional features incrementally as they become available. However, no current deep learning or FIL methods have been proposed to directly learn tabular data with missing values without requiring imputation.
Figure 1: No Imputation Incremental Learning (NIIL) algorithm uses $P$ fixed-sized overlapping feature partitions to incrementally train a Feature-Tokenized Transformer (FTT). Attention masks: $M_{1}$ operates as the attention column mask and $M_{2}$ as the attention row mask to exclude missing feature values from attention scoring.
@@
The state-of-the-art methods for handling missing values are broadly categorized into traditional and deep imputations. Traditional methods include MICE [12], missForest [13], and MLP in multiple imputations [5]). These methods use regression models to estimate missing values, assuming relationships between features. Deep imputation methods include Generative Adversarial Imputation Nets (GAIN) [14], diffusion-based imputation (Diffputer) [15], variational autoencoders [16], and denoising autoencoders [17]. Sun et al.  [18] report that GAIN has limited performance when missing values are at random (MAR) or not at random (MNAR), where MICE often outperforms GAIN. However, generative adversarial networks (GANs) can be unstable during optimization, particularly in terms of model convergence[19]. Diffusion-based methods use a simplified assumption about feature distributions (e.g., Gaussian distribution), which may not be optimal to yield the most accurate estimation of missing values. Chen et al. [20] report the superiority of GAIN over its diffusion counterparts, such as MissDiff [21] and CSDI_T [22]. These data-centric assumptions and limitations of the imputation methods explain why no single imputation model is ideal for all data sets. On the other hand, DL methods for missing value imputation have failed to attract domain applications due to high computational costs, lack of trust, and interpretability [23, 24]. In fact, several DL methods directly learn data with missing values by jointly imputing and classifying in a single framework [25, 26]. However, such an approach has the computational requirements of initializing and updating missing-value estimates with an imputation error. Joint learning of imputation and classification may bias the imputed values for given classification targets. Therefore, a DL framework capable of directly learning from observed values, excluding the contribution or imputation of missing values, would eliminate computational costs and biases due to synthetic data generation. In fact, decision tree-based ML models such as Random Forest and XGBoost [27] use “Missing Incorporated in Attributes” (MIA)[28] to directly learn data with missing values. However, traditional ML methods can be prone to overfitting high-dimensional feature space when incremental learning is not viable.
@@ 1.2 Motivation and contributions
This article makes five key contributions to address the challenges in learning tabular data with missing values. First, we present one of the first studies to streamline tabular data with missing values in a deep attention learning framework without requiring imputation or even initialization of missing values. Second, the proposed method derives and retrofits two attention masks to a feature- tokenized transformer to exclude missing entries from attention scoring. Third, we propose a new feature incremental learning strategy to enable efficient attention learning of O ($n^{2}$), which otherwise results in an out-of-memory problem in large feature spaces [32]. Fourth, the proposed incremental learning utilizes data in small and fixed-size feature partitions, rather than gradually increasing feature sizes. A fixed and small partition allows for more efficient learning of a large-dimensional feature space than conventional feature incremental learning methods. Fifth, the entire deep attention framework is designed to be computationally efficient and outperform state-of-the-art baseline methods, both with and without missing value imputations.
@@ 2.1 Preliminaries
2.2 Proposed NIIL algorithm
The proposed No Imputation Incremental Learning (NIIL) completely avoids imputation or initialization of missing values $X_{miss}$. Suppose a labeled tabular data set with missing values $D=\{X,y\}$ has $n$ samples and $d$ features. First, we sort the features $f_{1},f_{2},\dots,f_{m}$ in ascending order of missing rates. The sorted features are partitioned into subsets of $k$ features with $s=ceil(\frac{k}{2})$ overlaps, as shown in Figure 1. The overlap between two consecutive feature subsets facilitates incremental learning using $P$ partitions of features, as follows.
@@ 2.2 Proposed NIIL algorithm
These $P$ partitions are used to achieve incremental learning using a feature-tokenized transformer (FTT) [11, 33, 34] as follows. An FTT classifier model ($FTT_{1}$) is trained using the first partition $P_{1}=X\{f_{1},...,f_{k}\}$ of $k$ features. The next incremental learning session uses the second partition $P_{2}=X\{f_{s},...,f_{k+s}\}$ to train the previous $FTT_{1}$ model to $FTT_{2}$. Incremental learning continues until the $FTT_{P}$ model is obtained using the last feature partition, as illustrated in Figure 1. The trained $FTT_{P}$ model directly classifies a separate test data fold with missing values. Notably, FTT tokenizes features and uses a pre-trained language model to obtain embeddings for feature names and categorical features.In contrast, numerical values are passed through a linear projection layer to obtain corresponding embeddings. The tabular data representation is initialized by aggregating the embeddings of feature names and numerical and categorical feature values, which are subsequently fine-tuned for classification using a gated transformer. The proposed NIIL algorithm is presented in Algorithm 1. Missing values ($X^{miss}$) of X are excluded from attention scoring using attention masks designed as follows.
Algorithm 1 The NIIL Algorithm
@@ 3 Experiments
Table 1: Summary of the benchmark tabular data sets with heterogeneous data types and structures.
@@ 37 Diabetes 768 8 8 0 2
Naturally Missing 55 Hepatitis 155 19 6 13 2
@@
We simulate two types of missing values: 1) missing completely at random (MCAR) and 2) missing not at random (MNAR) at a missing rate between 10% and 50% at 10% intervals. Two data sets (IDs 55 and 6332) have naturally occurring missing values and are analyzed separately without simulating additional missing value types and rates.
@@ 3.1 Tabular data sets
Tabular data sets with large sample sizes (>>10,000) are often selected to facilitate deep learning [35, 36]. However, most tabular data sets in practice have relatively smaller sample sizes (<<1000) [37]. For generalizability, we have selected 15 diverse data sets with a variety of numerical and categorical feature mixtures from the OpenML [38] repository. The sample size of these data sets ranges from 155 to 2109, while the feature dimension is between 7 and 39. A summary of the tabular data sets used in this article is in Table 1, where the first 13 data sets are available in complete form without missing values. The remaining two data sets (IDs 55, and 6332) have missing values that occur naturally at the rate of 48.8% and 48.7%, respectively.
@@ 3.2 Model implementation and evaluation
The FTT classifier, adopted from [11], consists of a transformer with two encoder layers. The encoder layers transform numerical and categorical features into 128-dimensional embeddings. Each Transformer layer includes eight attention heads and a feedforward network with a 2048-dimensional hidden layer with ReLU activation functions and a dropout rate of 0.3. We have used the Adam optimizer with a fixed learning rate of $1\times 10^{-5}$ and no weight decay. The model is trained for up to 300 epochs using a batch size of 128. Early stopping is used with a patience of 50 epochs, terminating training when the validation loss ceases to improve for 50 consecutive epochs.
A five-fold cross-validation obtains the average area under the receiver operating characteristic curves (AUC) and the F1 scores. The average rank order of each method is reported in 15 tabular data sets based on AUC scores [8]. The five missing rates from 10% to 50% in 15 data sets yield 75 experimental scenarios for each missing value type. As in [39], we use a win matrix to compare methods in pairs by showing the percentage of experimental scenarios in which one method outperforms the other. The resilience of a model to varying missing rates is evaluated by comparing the AUC scores against the reference AUC score. The reference AUC score of a data set is obtained in its complete form before simulating missing values. We primarily evaluate the proposed incremental learning method for three partition sizes, k = 2, 3, and 4. Our analysis in Figure 5 suggests that FIL using feature partitions is computationally more efficient than learning the complete feature space ($d$ features) when k is less than $\lceil d/2\rceil$. Therefore, the performance of our model is reported for partitions when $k=\frac{d}{2}$.
@@ 3.3 Baseline methods
Our baseline methods have two types: one type requires imputation before classification, and the other can directly classify data without imputation. The requirement of imputation before classification is fulfilled by median value imputation, MICE [12], GAN-based imputation method GAIN [14], and diffusion-based imputation method, Diffputer [15]. Unlike mean values, median values are robust to outliers and are recommended for low missing rates. The subsequent classification of imputed data sets is achieved using two state-of-the-art models for tabular data, Gradient boosting trees (GBTs) and FTT [11]. On the other hand, XGBoost [27] and LightGBM [40] can use the Missing Incorporated in Attributes (MIA) strategy [28], as mentioned in Section 1.1, to directly classify data with missing values. XGBoost and LightGBM treat missing values as distinct categories during decision tree splits. The FTT classifier can directly classify missing values using an attention mask, presented in Equation 2.2. The proposed NIIL method is compared against 11 baseline methods: 1) median - GBT, 2) median - FTT, 3) MICE - GBT, 4) MICE - FTT, 5) GAIN - GBT, 6) GAIN - FTT, 7) Diffputer - GBT, 8) Diffputer - FTT, 9) AM - FTT, 10) MIA - XGBoost, and 11) MIA - LightGBM. All baseline methods are used with default hyperparameter settings.

### v5

@@ Abstract
Tabular data sets with varying missing values are prepared for machine learning using an arbitrary imputation strategy. Synthetic values generated by imputation models often raise concerns regarding data quality and the reliability of data-driven outcomes. To address these concerns, this article proposes an imputation-free incremental attention learning (IFIAL) method for tabular data with missing values. A pair of attention masks is derived and retrofitted to a transformer to directly streamline tabular data without imputing or initializing missing values. The proposed method incrementally learns partitions of overlapping and fixed-size feature sets to enhance the performance of the transformer. The average classification performance rank order across 17 diverse tabular data sets highlights the superiority of IFIAL over 11 state-of-the-art learning methods with or without missing value imputations. Additional experiments corroborate the robustness of IFIAL to varying types and proportions of missing data, demonstrating its superiority over methods that rely on explicit imputations. A feature partition size equal to one-half the original feature space yields the best trade-off between computational efficiency and predictive performance. IFIAL is one of the first solutions that enables deep attention models to learn directly from tabular data, eliminating the need to impute missing values. The source code for this paper is available at  [^1] https://anonymous.4open.science/r/ifial-435B.
@@ 1 Introduction
Missing values, occurring at varying rates and types, are ubiquitous in health sciences research using electronic health record (EHR) data. For two reasons, handling missing values remains critical to ensure unbiased data and reliable data-driven results. First, the state-of-the-art machine learning (ML) methods that promise precision medicine are not designed to learn data with missing values. Second, missing values are arbitrarily handled by excluding samples and features or imputing data with model-generated synthetic values to enable machine learning. Sample exclusion biases medical data with sicker patients when their healthy counterparts are known to have more missing values [hubbard2021]. Feature exclusion due to high missing rates can remove potentially important predictors. Furthermore, many health scientists are skeptical of using medical data imputed with synthetic values generated by models, especially when ML models are hard to interpret. Imputed values can alter feature importance, data statistics, and correlations between features, dampening the predictive ability of data [Arifeen2022]. For example, multiple imputations using chained equations (MICE) is arguably the most popular method in health science to impute missing values. MICE estimates missing values using a chain of linear regression models for individual features with two major limitations. First, regression models are ineffective when features have limited or no relationships with each other. Second, the relationship between features is rarely linear in practice. In other words, data statistics and missing value patterns should guide the modeling and selection of an imputation algorithm. Clearly, there is no best imputation method for all data sets or missing value types [payrovnaziri2021, Ribeiro2021, samad2022]. All concerns surrounding imputation methods underscore the need for alternatives to imputation in learning data with missing values. This paper proposes Imputation-free Incremental Attention Learning (IFIAL) of tabular data with missing values. IFIAL can directly classify tabular data with missing values, eliminating biases and concerns surrounding data exclusions and missing value imputation.
The remainder of this paper is organized as follows. Section 1.1 reviews the relevant literature on incremental feature learning and the learning of tabular data with or without missing values to motivate the contributions of our work. Section 2 introduces the proposed incremental learning and attention mask mechanism to develop the IFIAL framework. Section 3 discusses the experimental setup, data sets, missingness scenarios and baselines, and the evaluation method. Section 4 presents the performance of the proposed method and compares it with the state-of-the-art baselines, including robustness and ablation studies. Section 5 summarizes the key findings, provides insight into the results, and highlights the limitations of the proposed method. Finally, Section 6 concludes the paper.
@@ 1.1 Related work
The superiority of traditional ML in tabular data over deep learning is well documented in the literature [Borisov2022]. Feature incremental learning (FIL) is one of the cornerstones of deep learning (DL), which is not viable in traditional ML. Although highly successful with homogeneous image features [Li2022], incremental learning is challenging for tabular data due to a heterogeneous feature space [Liu2023]. Some recent work has proposed FIL on tabular data [Liu2023, Ahamed2024mambatab, Kim2024rail, Wang2024Transtab]. A common approach to FIL trains multiple models separately to accommodate new features and aggregate model predictions through a weighting mechanism [Kim2024rail, Liu2023]. Alternatively, mamba [Ahamed2024mambatab] and the transformer-based method [Wang2024Transtab] use a single model to train on additional features incrementally as they become available. However, the efficacy of FIL has not been investigated in learning tabular data with missing values, especially in handling missing values.
@@
The state-of-the-art methods for handling missing values are broadly categorized into traditional and deep imputations. Traditional methods include MICE [Lauren2021], missForest [Stekhoven2011], and MLP in multiple imputations [samad2022]). These methods train regression models to estimate and impute missing values, assuming the existence of relationships between features. Deep imputation methods include Generative Adversarial Imputation Nets (GAIN) [yoon2018gain], diffusion-based imputation (Diffputer) [diffputer], variational autoencoders [Hong2023], and denoising autoencoders [Shang2017]. Sun et al.  [sun2023deep] report that GAIN has limited performance when missing values are at random (MAR) or not at random (MNAR), where MICE often outperforms GAIN. However, generative adversarial networks (GANs) can be unstable during optimization, particularly in terms of model convergence [hyperimpute]. Diffusion-based methods employ a simplified assumption about feature distributions (e.g., a Gaussian distribution), which may not be optimal for yielding the most accurate estimation of missing values. Chen et al. [rethinking_diff] report the superiority of GAIN over its diffusion counterparts, such as MissDiff [ouyang2023missdiff] and CSDI_T [tabcsdi]. These data-centric assumptions and limitations of the imputation methods explain why no single imputation method is ideal for all data sets and missing value types. However, domain researchers have been reluctant to adopt DL methods for missing value imputation due to high computational costs, lack of trust, and interpretability [Bansal2021, Kamal2020]. Several DL methods have been proposed to learn data with missing values without using a standalone imputation step [Morvan]. However, similar methods first initialize missing values to subsequently learn a complete data representation during an end-to-end supervised learning task, such as classification and regression. Although the contribution of missing values to learning the complete data representation is masked out, these models do not distinguish between an observed zero value and zeros representing missing entries for a feature. Ideally, a DL framework should directly learn tabular data with missing values by streamlining only the observed values to address concerns about data quality and computational costs. In fact, decision tree-based ML models such as Random Forest and XGBoost [chen2016xgboost] use “Missing Incorporated in Attributes” (MIA)[twala2008mia] to directly learn data with missing values. However, traditional ML methods can be prone to overfitting high-dimensional feature spaces when incremental feature learning is not viable.
Figure 1: Three strategies for handling missing values in tabular data. Top: Standalone imputation of missing values before classification. Middle: Joint learning of complete data representation and classification. Bottom: (Proposed) - no imputation or initialization of missing values needed to streamline only the observed values for classification.
@@ 1.2 Motivation and contributions
  • We present one of the first studies to streamline tabular data with missing values in a deep attention learning framework without requiring imputation or even initialization of missing values.
  • The proposed method derives and retrofits two attention masks to a feature-tokenized transformer to exclude missing entries from attention scoring.
  • We propose a new incremental learning strategy to mitigate the effect of varying missing rates in the feature space and allow efficient attention learning of O ($n^{2}$), which otherwise results in an out-of-memory problem in large feature spaces [rabbani2024].
  • The proposed incremental learning partitions the data into small and fixed-size feature spaces, rather than gradually increasing feature sizes, allowing more efficient learning of a large-dimensional feature space.
  • The entire deep attention framework is designed to be computationally efficient and performancewise superior to state-of-the-art baselines, both with and without missing value imputations.
@@ 2.1 Preliminaries
Figure 2: Imputation-free Incremental Attention Learning (IFIAL) algorithm uses $P$ fixed-sized overlapping feature partitions to train a Feature-Tokenized Transformer (FTT) incrementally. Attention masks: $M_{1}$ operates as the attention column mask and $M_{2}$ as the attention row mask to exclude missing feature values from attention scoring.
2.2 Proposed IFIAL algorithm
The proposed Imputation-free Incremental Attention Learning (IFIAL) completely avoids imputation or initialization of missing values $X_{miss}$. Suppose a labeled tabular data set with missing values $D=\{X,y\}$ has $n$ samples and $d$ features. First, we sort the features $f_{1},f_{2},\dots,f_{m}$ in ascending order of missing rates. The sorted features are partitioned into subsets of $k$ features with $s=ceil(\frac{k}{2})$ overlaps, as shown in Figure 2. The overlap between two consecutive feature subsets facilitates incremental learning using $P$ partitions of features, as follows.
@@ 2.2 Proposed IFIAL algorithm
These $P$ partitions are used to achieve incremental learning using a feature-tokenized transformer (FTT) [Wang2024Transtab, huang2020, schambach2023] as follows. An FTT classifier model ($FTT_{1}$) is trained using the first partition $P_{1}=X\{f_{1},...,f_{k}\}$ of $k$ features. The next incremental learning session uses the second partition $P_{2}=X\{f_{s},...,f_{k+s}\}$ to train the previous $FTT_{1}$ model to $FTT_{2}$. Incremental learning continues until the $FTT_{P}$ model is obtained using the last feature partition, as illustrated in Figure 2. The trained $FTT_{P}$ model directly classifies a separate test data fold with missing values. Notably, FTT tokenizes features and uses a pre-trained language model to obtain embeddings for feature names and categorical features. In contrast, numerical values are passed through a linear projection layer to obtain corresponding embeddings. The tabular data representation is initialized by aggregating the embeddings of feature names and numerical and categorical feature values, which are subsequently fine-tuned for classification using a gated transformer. The proposed IFIAL algorithm is presented in Algorithm 1. Missing values ($X^{miss}$) of X are excluded from attention scoring using attention masks designed as follows.
Algorithm 1 The IFIAL Algorithm
@@ 3 Experiments
Table 1: Summary of tabular data sets with heterogeneous data types and structures used for benchmarking.
@@ 37 Diabetes 768 8 8 0 2
Natural Missing Values 55 Hepatitis 155 19 6 13 2
@@
41162 Kick 72983 10 4 6 2
41440 Okcupid-stem 50789 10 9 1 3
We perform experiments on three types of missing values: 1) missing completely at random (MCAR), 2) missing not at random (MNAR), and 3) missing values that appear naturally with the data. Here, MCAR and MNAR types are simulated at a missing rate ranging from 10% to 50% in increments of 10%.
@@ 3.1 Tabular data sets
Table 1 summarizes the tabular data sets used for the benchmarking. For generalizability, we have selected 17 diverse data sets of varying sizes and a mix of numerical and categorical features from the OpenML repository [OpenMLCC18]. Notably, tabular data sets with large sample sizes (>>10,000) are often selected to facilitate deep learning [caruso2024, FTT_Gorishniy2021]. However, most tabular data sets in practice have relatively smaller sample sizes (<<1000) [rabbani-g-ceals-2024], which are more challenging data sets for deep learning methods. Therefore, our selected data sets have sample sizes ranging from 155 to 72983, while the feature dimension ranges from 7 to 39. The first 13 data sets are available in their complete form, without missing values, and are used to simulate the MCAR and MNAR types of missing values following [samad2022]. The other four data sets with IDs 55, 6332, 41162 and 41440 have natural missing values at the rate of 48.8%, 48.7%, 30.2%, and 20.6%, respectively.
@@ 3.2 Model implementation and evaluation
The FTT classifier, adopted from [Wang2024Transtab], consists of a transformer with two encoder layers. The encoder layers transform numerical and categorical features into a unified 128-dimensional embedding. Each Transformer layer includes eight attention heads and a feedforward network with a 2048-dimensional hidden layer, ReLU activation, and a dropout rate of 0.3. We have used the Adam optimizer with a fixed learning rate of $1\times 10^{-5}$ and no weight decay. The model is trained for up to 300 epochs using a batch size of 128. Early stopping is used with a patience of 50 epochs, terminating training when the validation loss stops improving for 50 consecutive epochs. The proposed and all baseline models are validated using Optuna [optuna_akiba]. For each experiment, we run 100 Optuna trials, randomly sampling different hyperparameter settings and evaluating them on the validation split. The model achieving the best validation performance is then selected and evaluated on the corresponding test fold. The hyperparameter search spaces are reported in Table 6.
A five-fold cross-validation yields the average area under the receiver operating characteristic (ROC) curve (AUC). The average rank order of each method is reported in 13 tabular data sets based on AUC scores [Borisov2022]. The five missing rates from 10% to 50% in 13 data sets yield 65 experimental scenarios for each missing value type. As in [tokar2024icet], we use a win matrix to compare methods in pairs, showing the percentage of experimental scenarios in which one method outperforms the other. The resilience of a model to varying missing rates is evaluated by comparing the AUC scores against the reference AUC score. The reference AUC score of a data set is obtained in its complete form before simulating missing values. We primarily evaluate the proposed incremental learning method for three partition sizes, k = 2, 3, and 4. Our analysis in Figure 6 suggests that FIL using feature partitions is computationally more efficient than learning the complete feature space ($d$ features) when k is less than $\lceil d/2\rceil$. Therefore, we report the performance of our model for partitions with $k=\frac{d}{2}$.
@@ 3.3 Baseline methods
Our baseline methods can be categorized into two types: one requires imputation before classification, and the other can classify data directly without a standalone imputation. The imputation requirement before classification is satisfied by median value imputation, MICE [Lauren2021], GAN-based imputation method GAIN [yoon2018gain], and diffusion-based imputation method, Diffputer [diffputer]. Unlike mean values, median values are robust to outliers and are recommended for low missing rates. The classification of imputed data sets is achieved using two state-of-the-art models for tabular data, Gradient boosting trees (GBTs) and FTT [Wang2024Transtab]. On the other hand, XGBoost [chen2016xgboost] and LightGBM [ke2017lightgbm] can use the Missing Incorporated in Attributes (MIA) strategy [twala2008mia], as mentioned in Section 1.1, to directly classify data with missing values. XGBoost and LightGBM treat missing values as distinct categories during decision tree splits. The FTT classifier can directly classify missing values using an attention mask, presented in Equation 2.2. Therefore, an AM-FTT framework, similar to  [caruso2024], is considered a baseline. The proposed IFIAL method is compared against 11 baseline methods: 1) median - GBT, 2) median - FTT, 3) MICE - GBT, 4) MICE - FTT, 5) GAIN - GBT, 6) GAIN - FTT, 7) Diffputer - GBT, 8) Diffputer - FTT, 9) AM - FTT, 10) MIA - XGBoost, and 11) MIA - LightGBM.